OpenStudy (anonymous):

Find the sum of the geometric sequence. 3/2, 3/8, 3/32, 3/128, 3/512

ganeshie8 (ganeshie8):

use the partial sum formula

ganeshie8 (ganeshie8):

\(\large S_n = a\left(\frac{1-r^n}{1-r}\right)\)

ganeshie8 (ganeshie8):

figure out the numbers and plugin

OpenStudy (anonymous):

ummmm can you give me a hint? Is the fraction going to be big or small?

ganeshie8 (ganeshie8):

first term, a = 3/2

ganeshie8 (ganeshie8):

common ratio, r = ?

OpenStudy (anonymous):

r = 3?

OpenStudy (anonymous):

Thanks for the message. I get it now.

ganeshie8 (ganeshie8):

nope r = ratio of adjacent terms r = ( 3/8 ) / ( 3/2 ) = 1/4

ganeshie8 (ganeshie8):

count the number of terms, n = ?